# Prediction Instability in Machine Learning Ensembles

### v1

@@ 3 An Ensemble Choice Discontinuity Theorem
In this section we will establish our main result. We will argue that there is a fundamental tradeoff between information use and stability of ensemble predictions and choices in the sense that no aggregation procedure satisfies defintions 2.1, 2.2, 2.3, and 2.4, a result that we refer to in shorthand as ensemble choice discontinuity (ECD). Our focus here will be on building intuition for why our ECD theorem is true via sketching the arguments for it. The following section will expand on implications with examples and discussion. Full proofs for all results are in the appendix.
Central to the proof of our ECD theorem is the notion of transitivity:
Definition 3.1.
Say that $C(\cdot,s)$ satisfies transitivity if for all $y,y^{\prime},y^{\prime\prime}\in Y$ and all $s\in S(Y)$ we have that $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$ implies $y\in C(\{y,y^{\prime\prime}\},s)$.
@@
In the first step of our argument we show that any ensemble choice aggregator that satisfies insertion/deletion consistency will be transitive, formalized in this Lemma:
Lemma 3.2.
$C(\cdot,s)$ satisfies insertion/deletion consistency if and only if it is transitive.
@@
So, given an ensemble choice aggregator that is insertion/deletion consistent, and so transitive, the next step is to argue that transitivity implies that at least one information use requirement must be violated. The strategy to show this follows insights from Geanakoplos [2005] and Fey [2014]. The first step is to use ensemble unanimity to identify a set of scores in which there is a critical model – a candidate for a model in the ensemble whose scores might completely determine ensemble choices. The second part of the argument is to apply ensemble unanimity, transitivity, and model choice reversal to a series of scores derived from making “small” changes to the original set of scores used to identify the critical model. In particular, where model choice reversal holds, if no model reverses its scores for one label versus another then the ensemble cannot change its choices between the two labels. Formally:
Lemma 3.3.
@@
Suppose that $C(\cdot,s)$ respects model choice reversal. Then $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ iff $s_{j}^{\prime}(\cdot,y)>s_{j}^{\prime}(\cdot,y^{\prime})$ for all $j$ implies that $C(\{y,y^{\prime}\},s)=C(\{y,y^{\prime}\},s^{\prime})$.
@@
Lemma 3.3 allows us to show that each of the series of model scores must yield the same choice as a set of scores in which the critical model determines the choice; we conclude that $C(\cdot,s)$ must not be a nondegenerate ensemble choice aggregator, leading to our main result, which we now state:
Theorem (Ensemble Choice Discontinuity).
Suppose that $|Y|\geq 3$. If $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator then it must violate at least one of insertion/deletion consistency, ensemble unanimity, or model choice reversal.
4 Implications
What does this theoretical result mean for how we should expect machine learning ensembles to behave in practical settings and what lessons do we take away? In general, the ECD theorem tells us that if we would like to have a machine learning ensemble use the insights of multiple models in making predictions and choices (definition 2.2) then we must accept the risk that the ensemble makes those predictions and choices in a counterintuitive or unstable fashion at least some of the time, ignoring model agreement (definition 2.3), changing its mind in a way that the underlying models cannot justify (definition 2.4) or in response to the presence of label options it would not predict or choose (definition 2.1).
Violations of insertion/deletion consistency are a form of discontinuity in the scoring space in the sense that changes to the unchosen portion of the set of label values can induce changes in the ensemble prediction or choice even if the individual model scores are held constant. Violations of model choice reversal are a form of discontinuity in the underlying information in the sense that two scores can both be consistent with the same ordering of the set of label values for all models and yet result in different ensemble predictions or choices.
We expand on this with some examples. To make this discussion concrete, a natural way to use an ensemble to make decisions about labels would be to aggregate the scores of individual models and then require that the ensemble choose the label that has the highest aggregated score. This method of making choices is clearly insertion/deletion consistent and transitive under any method of score aggregation.
In a voting ensemble individual model scores are combined to make a final prediction or decision through either hard or soft voting. In hard voting the voting ensemble predicts the class label that receives the plurality of votes from the constituent models in which case the ensemble choice aggregator will be:
$C(Y^{*},s)$
$\hskip 8.53581pt=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\sum_{j}I\left(y\in\operatornamewithlimits{argmax}\left\{s_{j}(\cdot,y^{\prime})\right\}\right)\right\}.$ (1)
In soft voting the ensemble averages the scores assigned to each response label value by the models in which case the ensemble choice aggregator will be:
$C(Y^{*},s)=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\frac{1}{m}\sum_{j=1}^{m}s_{j}(\cdot,y)\right\}.$ (2)
In either case $C(\cdot,s)$ is nondegenerate and satisfies unanimity, and so the ECD theorem implies a violation of model choice reversal in multiclass or regression problems. Relating this back to our random forest motivating example, the use of a unit vote for the most popular class [e.g. Breiman, 2001] implemented as a form of soft voting implies a violation of model choice reversal; the decision making inconsistency that we used to motivate our argument is indeed ubiquitous.
In stacked generalization we train a collection of models on a data set and then learn an aggregation method that assembles the output of these models into a single prediction as a secondary learning task [Wolpert, 1992]. One approach to this would be to learn a set of weights that could be used to aggregate the individual models. For example, in the case where we have a set of Bayesian regression models combined via learned weights the posterior distribution over the parameters of each model induces a score for each $y\in Y$ when given $x$ in the form of the model evidence $p_{j}(y|x)$. Thus the model scoring function is $s_{j}(x,y)=p_{j}(y|x)$ and the ensemble choice aggregator is:
$C(Y^{*},s)=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\sum_{j}w_{j}s_{j}(x,y)\right\}$
where $w_{j}$ is the learned weight for model $j$. If $w_{j}<0$ for some $j$ then the ensemble choice aggregator will violate ensemble unanimity. If $w_{j}=0$ for all but one model then this is not a nondegenerate ensemble choice aggregator, for example as a consequence of using a model specification method, e.g. the adaptive lasso [see Zou, 2006]. If $w_{j}>0$ for all $j$ then it is a nondegenerate ensemble choice aggregator and satisfies ensemble unanimity, and so by the ECD theorem is guaranteed to violate model choice reversal. Learning simply imposes a particular violation as an accidental by-product of optimizing the loss function.
Finally, reinforcement learning methods have been introduced in which multiple agents work together in smart collaboration to maximize rewards via voting [e.g. Xu et al., 2019]. As a simple example, in $Q$-learning, agents learn to encode the long-term value for taking actions a function mapping from an action $a$ and state of the world $\sigma$, the so-called $Q$-value, $Q(\sigma,a)$. If each agent learns its own $Q$-value, these values are aggregated via soft voting, and the action associated with the largest is chosen then the ECD theorem implies a violation of model choice reversal; similar states of the world with very similar state-action values will sometimes lead to radically different decisions.
We summarize these observations in the following corollary of the ECD theorem:
Corollary 4.1.
If $C(\cdot,s)$ satisfies equations 1 or 2 then it violates model choice reversal.
Discussion
Specifying a choice aggregator as a nondegenerate ensemble and satisfying ensemble unanimity are otherwise relatively undemanding requirements and it is easy to write down score aggregators that will use the input of more than one model and respect agreement among all models; indeed a simple weighted average where all weights are positive works. Model choice reversal–the requirement that the ensemble cannot change its mind unless some of the underlying models do so–is more difficult to guarantee, and so we focus discussion on this property.
Intuitively, model choice reversal is a form of smoothnesss. When model choice reversal holds “small” changes in individual model scores cannot provoke “large” changes in aggregated decisions. On the other hand when model choice reversal is violated the implication is that the ensemble cannot guarantee that 1) small changes to model scoring functions via training updates or 2) two nearly identical inputs will generate similar decisions. Small changes to the scores will generate large changes to the outputs in at least some part of the scoring space; thus aggregation of multiple models must necessarily introduce “discontinuities” in the decisions made by an ensemble. Inability to guarantee equal treatment of nearly equivalent cases seems to violate basic fairness requirements. We suspect that this will manifest as unpredictable and unintuitive behavior that complicates the explainability of an ensemble and compromises the user experience for a non-expert.
For example, model choice reversal is closely related to the idea of monotonicity:
Definition 4.2.
Say that $C(\cdot,s)$ is monotonic if for any $y,y^{\prime}\in Y$ and any $s,s^{\prime}\in S(Y)$ such that $s_{j}(\cdot,y)\geq s_{j}(\cdot,y^{\prime})$ implies $s^{\prime}_{j}(\cdot,y)\geq s_{j}^{\prime}(\cdot,y^{\prime})$ for all $j$ we have that $y\in C(\{y,y^{\prime}\},s)$ implies that $y\in C(\{y,y^{\prime}\},s^{\prime})$.
A monotonic aggregator is one in which more individual model support for a particular label will not make the ensemble less likely to choose it. It is straightforward to show that any monotonic ensemble score aggregator will respect model choice reversal (see Lemma 5.2 in the appendix). The following corollary is then an immediate consequence of our ECD theorem:
Corollary 4.3.
When $|Y|\geq 3$ then if $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator that satisfies insertion/deletion consistency and ensemble unanimity it violates monotonicity.
So, where model choice reversal is violated monotonicity will also be violated and there will be sets of scores–and inputs associated with them–where more individual model support for a label will rule that label out as an ensemble choice, an intrinsic property of an ensemble that downstream users may find difficult to grasp and so one that undermines their trust.
The inconsistency of model choice reversal with the other properties is also tied to the capacity of the individual models within the ensemble to represent the relationship between features and labels. We observe that in our setting we can think about formalizing model representational capacity in terms of two properties. First, the range of the scoring function, $S_{j}(Y)$, bounds the ability of the model to encode information about the label. When we defined this range above we required that it include scores consistent with any ordering of the label values. Second, if each of these scores is mapped to from some point in the feature space–if $s_{j}(\cdot)$ is surjective–then we can conclude that the model has enough capacity to represent any relationship between features and labels.
This situation is sufficient to guarantee that violations of the properties in our ECD theorem are “observable” in the sense that they can actually be encountered at inference because real points in feature space lead to the scores associated with violations of definitions 2.2, 2.3, and 2.4. Restating this discussion formally:
Corollary 4.4.
@@ Discussion
Suppose that $|Y|\geq 3$ and that $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator. If $s(\cdot)$ is surjective then there exists $X^{*}\subseteq X$ such that for all $x^{*}\in X^{*}$ we have that $C(\cdot,s(x^{*}))$ violates at least one of insertion/deletion consistency, ensemble unanimity, or model choice reversal.
High capacity individual models increase the risk of violations of at least one of the properties analyzed here and–to the extent that the other properties are easily satisfied–to violations of model choice reversal. It follows that there is a fundamental tradeoff between the flexibility of the individual models in an ensemble and violations of model choice reversal.
This also suggests that restrictions to the ranges of individual models or to the mapping from feature space to scores may eliminate portions of the scoring space that violate model choice reversal. For example, in the weighted averaging of Bayesian models discussed above if the distributional assumptions of the individual Bayesian regressions impose a unimodal structure on the model evidence this would amount to a restriction on the range of the individual models and so on the domain of the aggregator; under this restricted domain using majority rule as an ensemble choice aggregator would satisfy our three properties [see Black, 1948]. It also strikes us that the requirements on the model scoring function range and the surjectivity of $s_{j}(\cdot)$ for Corollary 4.4 are likely much stronger assumptions in regression problems than in classification problems; though the ECD theorem holds in a technical sense in regression we speculate that the risk of running into model choice reversal violations is lower in regression problems.
5 Conclusion

### v5

@@ III An Ensemble Aggregated–Prediction Instability Theorem
In this section we will argue that there is a fundamental tradeoff between information use and stability of ensemble predictions and choices in the sense that no aggregation procedure satisfies defintions II.4, II.5, II.6, and II.7. Our focus here will be on building intuition for why our theorem is true via sketching the arguments for it. The following section will expand on implications with examples and discussion. Full proofs for all results are in the appendix.
Transitivity plays a central role in the proof of our result:
Definition III.1.
Say that $C(\cdot)$ satisfies transitivity if, for all $y,y^{\prime},y^{\prime\prime}\in Y$ and any $z\in S(Y)$, we have that $y\in C(\{y,y^{\prime}\},z)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},z)$ implies $y\in C(\{y,y^{\prime\prime}\},z)$.
@@
We being by assuming that all of our conditions hold. In the first step of our argument we show that any ensemble choice aggregator that satisfies insertion/deletion consistency will be transitive, formalized in this Lemma:
Lemma III.2.
$C(\cdot)$ satisfies insertion/deletion consistency if and only if it is transitive.
@@
So, given an ensemble choice aggregator that is insertion/deletion consistent, and thus transitive, the next step is to argue that transitivity implies that at least one information use requirement must be violated. The strategy to show this follows insights from Geanakoplos [14] and Fey [10]. We use ensemble unanimity to identify a set of scores in which there is a critical model – a candidate for a model in the ensemble whose scores might completely determine ensemble choices. The second part of the argument is to apply ensemble unanimity, transitivity, and model choice reversal to a series of scores derived from making “small” changes to the original set of scores used to identify the critical model. In particular, where model choice reversal holds, if no model reverses its scores for one label versus another then the ensemble cannot change its choices between the two labels. Formally:
Lemma III.3.
@@
Suppose that $C(\cdot)$ respects model choice reversal. Then $z_{j,d(y)}>z_{j,d(y^{\prime})}$ if and only if $z^{\prime}_{j,d(y)}>z^{\prime}_{j,d(y^{\prime})}$ for all $j$ implies that $C(\{y,y^{\prime}\},z)=C(\{y,y^{\prime}\},z^{\prime})$.
@@
Lemma III.3 allows us to show that each of the series of model scores must yield the same choice as a set of scores in which the critical model determines the choice; we conclude that $C(\cdot)$ must not be a nondegenerate ensemble choice aggregator, a contradiction that establishes that there must be a set of scores on which at least one of our definitions is violated. Formally:
Theorem III.4.
(Aggregated–Prediction Instability) Suppose that $|Y|\geq 3$, that all individual models are unconstrained learners, and that $C(\cdot)$ is a non-degenerate ensemble choice aggregator on $S(Y)$. Then:
  1. 1.
@@
    There exists $S^{*}\subseteq S(Y)$ such that $C(\cdot)$ violates at least one of insertion/deletion consistency, ensemble unanimity, or model choice reversal on $S^{*}$.
  2. 2.
    If $s(x)$ is surjective then $S^{*}$ is observable by $s(x)$.
IV Implications and Discussion
Next we turn to laying out what this theoretical result means for how we should expect machine learning ensembles to behave in practical settings and what lessons we take away.
IV-A Aggregation creates problems
In general, theorem III.4 tells us that if we would like to have a machine learning ensemble use the insights of multiple models in making predictions and choices (definition II.5) then we must accept the risk that the ensemble makes those predictions and choices in a counterintuitive or unstable fashion at least some of the time, ignoring model agreement (definition II.6), changing its mind in a way that the underlying models cannot justify (definition II.7) or in response to the presence of label options it would not predict or choose (definition II.4).
Failures of these properties are forms of “instability” in the sense that they generate sudden changes in the output of the ensemble. Violations of insertion/deletion consistency are a form of discontinuity in the scoring space in the sense that changes to the unchosen portion of the set of response labels can induce changes in the ensemble prediction or choice even if the individual model scores are held constant. Violations of ensemble unanimity are also a form of discontinuity in the scoring space in that more model agreement on the ordering of the response labels can lead to the ensemble ignoring labels which all agree to be more or most likely. Violations of model choice reversal are a form of discontinuity in the underlying information in the sense that two scores can both be consistent with the same ordering of the set of label values for all models and yet result in different ensemble predictions or choices.
Finally, we note that part 1) of theorem III.4 implies that violations of these properties are intrinsic characteristics of the method by which the individual models in the ensemble are aggregated, $C(\cdot)$, and do not depend on what $s(x)$ has learned. Choosing a particular aggregation method determines which properties will be violated and where in the scoring space those violations will occur.
IV-B Capacity of individual models makes problems observable
Violations of these properties–and particularly the extent to which they will actually be encountered by a user during model inference–are also tied to the capacity of the individual models within the ensemble to represent the relationship between features and labels. In our setting we formalized model capacity in terms of two properties: unconstrained learners and observability. First, the range of the scoring function, $S_{j}(Y)$, bounds the ability of the model to encode information about the label. When the models in the ensemble are unconstrained learners this range includes scores consistent with all orderings of the response labels, so that any ordering can be learned. This is sufficient to guarantee that there will be learnable regions of the scoring space where aggregation will lead to violations of our properties.
Second, if each vector of scores in $S(Y)$ is observable in the sense that it is mapped to from some point in the feature space then we can conclude that the model has enough capacity to represent any relationship between features and response labels. If $s(x)$ is surjective this is sufficient to guarantee that all scores in $S(Y)$ are observable and could be encountered during inference. Consequently, it is also sufficient to guarantee that violations of our properties are observable in the sense that they can be encountered at inference because real points in feature space lead to the scores associated with those violations.
High capacity individual models increase the risk of observable violations of at least one of the properties analyzed here. In this sense, theorem III.4 provides an additional axiomatic rationale for the use of weak learners–limited capacity individual models–within an ensemble.
IV-C Popular aggregation procedures violate model choice reversal
Specifying the aggregation method as a non-degenerate ensemble choice aggregator that satisfies ensemble unanimity is relatively undemanding and it is easy to write down score aggregators that will use the input of more than one model and respect agreement among all models; indeed a simple weighted average where all weights are positive works. Model choice reversal–the requirement that the ensemble cannot change its mind unless some of the underlying models do so–is more difficult to guarantee, and multiple popular approaches to ensembling will violate it.
We expand on this with some examples. A natural, perhaps even ubiquitous, way to use an ensemble to make decisions about response labels would be to aggregate the scores of individual models and then require that the ensemble choose the label that has the highest aggregated score. This is clearly insertion/deletion consistent and transitive under any method of score aggregation.
In a voting ensemble individual model scores are combined to make a final prediction or decision through either hard or soft voting. In hard voting the voting ensemble predicts the class label that receives the plurality of votes from the constituent models in which case:
$C(Y^{*},s(x))$
$\hskip 8.53581pt=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\sum_{j}I\left(y\in\operatornamewithlimits{argmax}\left\{s_{j}(x,y)\right\}\right)\right\}.$ (1)
In soft voting the ensemble averages the scores assigned to each response label value by the models in which case:
$C(Y^{*},s(x))=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\frac{1}{m}\sum_{j=1}^{m}s_{j}(x,y)\right\}.$ (2)
Either way, $C(\cdot)$ is a nondegenerate ensemble choice aggregator and satisfies unanimity, and so theorem III.4 implies a violation of model choice reversal in multiclass or regression problems.
These voting procedures are commonly used to aggregate individual models. For example, tree ensembles, including those generated by e.g. random forest, XGBoost or Light GBM, often apply soft voting or weighted soft voting for aggregation. Even reinforcement learning methods can apply soft voting; in [e.g. 33] multiple agents work together in smart collaboration to maximize rewards via voting. As a simple example, in $Q$-learning, agents learn to encode the long-term value for taking actions in a function mapping from an action $a$ and state of the world $\sigma$, the so-called $Q$-value, $Q(\sigma,a)$. If each agent learns its own $Q$-value, these values are aggregated via soft voting, and the action associated with the largest is chosen then:
$C(Y^{*},Q(\sigma))=\operatornamewithlimits{argmax}_{a\in Y^{*}}\left\{\frac{1}{m}\sum_{j=1}^{m}Q_{j}(\sigma,a)\right\}.$ (3)
We summarize these observations in the following corollary:
Corollary IV.1.
If $C(\cdot)$ satisfies equations 1, 2, or 3 then there exists $S^{*}\subseteq S(Y)$ such that it violates model choice reversal on $S^{*}$.
The use of a unit vote for the most popular response label or action, [e.g. 6], implemented as a form of soft voting implies a violation of model choice reversal. The decision making inconsistency that we used to motivate our argument is indeed ubiquitous.
IV-D Violations of model choice reversal complicate explainability
As we discussed above, model choice reversal is a form of smoothness. When model choice reversal holds “small” changes in individual model scores–no modeling changing its mind–cannot provoke “large” changes in aggregated decisions. On the other hand when model choice reversal is violated small changes to individual model scores will generate large changes to the outputs in at least some part of the scoring space. Small changes to individual model scores can come about in multiple ways, but two obvious pathways are 1) through training updates to the models and 2) through conducting inference on different points in feature space. In the case of training updates, violations to model choice reversal mean that the ensemble cannot guarantee similar predictions or choices by the ensemble before and after the training update even if that update yields very minimal change to the models. In the case of inference for different points in the feature space the ensemble cannot guarantee that two nearly identical points in feature space will generate similar decisions even when the individual models are locally smooth. We suspect that this will manifest as unpredictable and unintuitive behavior that complicates the explainability of an ensemble and compromises the user experience for a non-expert.
IV-E Ensembles cannot learn their way out
Learning the ensemble choice aggregator along with the individual models will simply impose a particular violation as an accidental by-product of optimizing the loss function. As an example of this, in stacked generalization we train a collection of models on a data set and then learn an aggregation method that assembles the output of these models into a single prediction as a secondary learning task [32]. One approach to this would be to learn a set of weights that could be used to aggregate the individual models. For example, in the case where we have a set of Bayesian regression models combined via learned weights the posterior distribution over the parameters of each model induces a score for each $y\in Y$ when given $x$ in the form of the model evidence and the ensemble choice aggregator is:
$C(Y^{*},s(x))=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\sum_{j}w_{j}s_{j}(x,y)\right\}$
where $w_{j}$ is the learned weight for model $j$. If $w_{j}<0$ for some $j$ then the ensemble choice aggregator will violate ensemble unanimity. If $w_{j}=0$ for all but one model then this is not a nondegenerate ensemble choice aggregator, for example as a consequence of using a model specification method, e.g. the adaptive lasso [see 34]. If $w_{j}>0$ for all $j$ then it is a nondegenerate ensemble choice aggregator and satisfies ensemble unanimity, and so by theorem III.4 is guaranteed to violate model choice reversal.
V Asymptotic Limitations on Ensemble Aggregated–Prediction Instability
The discussion of the link between model capacity and ensemble aggregation instability highlights the role played by having surjective, unconstrained learners in creating violations of our aggregation prediction/choice properties. In turn, this suggests that restrictions to the ranges of individual models or to the mapping from feature space to scores may eliminate portions of the scoring space that violate our properties, and so promote stability in ensemble decision making. Following this reasoning, we conjecture that as the data available for training grows arbitrarily large this will act as an implicit restriction on the diversity of the models in the ensemble and thus the ranges and mappings of individual models, and so minimize the risk of prediction instability.
To formalize this intuition, we draw upon the notion of machine learning consistency. We suppose that there is a true conditional distribution over $Y$, say $p(Y=y|X=x)$, that reflects the probability of $y\in Y$ given full knowledge of the underlying data population (rather than the sample available for training). Given this true conditional distribution, it is well known [e.g. 3] that the Bayes classifier minimizes the risk of making incorrect predictions or choices, and so consistency is typically defined in a machine learning context as convergence in probability to the prediction or choice made by the Bayes classifier as the size of the data sample available for training grows.
Our version of consistency is similar, if slightly stronger:
Definition V.1.
Say that a sequence of model scoring functions $s_{j}(x|D_{n})$ is fully consistent if for all $\varepsilon>0$
00 0 $=\lim_{n\to\infty}p\left(\sum_{y,y^{\prime}}\int\limits_{X}\begin{array}[]{l}I\{s_{j}(x,y|D_{n})\leq s_{j}(x,y^{\prime}|D_{n})\}\\ \times I\{p(y|x)>p(y^{\prime}|x)\}p(x)dx\end{array}\geq\varepsilon\right).$
That is, a model scoring function is fully consistent if the probability that it disagrees with the pairwise ordering of the true conditional distribution is zero in the limit of arbitrarily large amounts of training data. [^8] Note that this is stronger than the usual definition of machine learning consistency. The usual definition only requires agreement between the model and Bayes classifier on the choice from the full set $Y$. Full consistency is sufficient but not necessary for this in the sense that if $s_{j}(x|D_{n})$ is fully consistent then a classifier defined on it in the usual way will converge to the Bayes classifier.
If all models in an ensemble are fully consistent then we have the following result:
Theorem V.2.
Suppose that $s_{j}(x|D_{n})$ is fully consistent for all $j$. Then for almost all classification problems as $n\to\infty$ there exists a non-degenerate ensemble choice aggregator $C(\cdot)$ that satisfies insertion/deletion consistency, ensemble unanimity, and model choice reversal for all $x\in X$ with probability one.
Consistency of the individual models in the ensemble means that asymptotically it is possible to find ensemble aggregation procedures that satisfy definitions II.4, II.5, II.6, and II.7.999The “almost all” in the statement of the theorem refers to the true conditional distribution – for any finite set $Y$ the set of distributions with ties between the elements of $Y$ is measure zero relative to the full set of distributions over $Y$. Also, although we have proved this result with asymptotic convergence of the models in the ensemble to the true conditional distribution in mind, agreement of the models in the ensemble with the order derived from the true conditional distribution is not necessary to eliminate prediction or choice instability. Rather, the key is simply agreement of the models with one another full stop, whatever distribution they ultimately agree on. Specifically, soft voting, which suffers from violations to model choice reversal in finite samples, will satisfy all properties with consistent models operating in asymptotic conditions. Prediction or choice instability is in part a finite sample issue.
Finally, it is worth noting that full consistency at the level of the ensemble is not strong enough; a consistent ensemble may still exhibit prediction or choice instability even in asymptotic conditions. To find ensemble choice aggregators that will not violate our properties we need all models in the ensemble to be fully consistent. Put another way, diversity among the models within the ensemble, the very thing that may be desirable for generalization error minimization, creates the prediction or choice instability we analyze here.
VI Conclusion